OpenStudy (anonymous):

what is the simplified form of (2x^6) (3x^1/2)? I got 6x^3 but it was marked wrong

OpenStudy (raden):

2 * 3 = 6 but x^6 * x^1/2 = x^(6 + 1/2) = x^(6.5) = x^13/2